1/2

William Blake,  William Blake,  William Blake,
The Tyger  Le tigre  Der Tiger
translated by Félix Rose  translated by Walter A. Aue

Tyger Tyger, burning bright, In the forests of the night; What immortal hand or eye, Could frame thy fearful symme-

try?

Tigre, tigre flamboyant

d’ardeur Dans les forêts de la nuit; Quel pouvoir immortel, inouï, Forma ta parfaite terreur?

Argiope information:

The view of this version is re-

stricted. Presumably there are copy-

right issues.

In what distant deeps or skies Burnt the fire of thine eyes? On what wings dare he aspire? What the hand, dare sieze the

fire?

Dans quels cieux, quelles pro-

fondeurs Brûlait le feu de ton regard? Quel vol osa donc vouloir? Et quelle main saisit ce feu?

And what shoulder and what

art, Could twist the sinews of thy

heart? And when thy heart began to

beat, What dread hand? and what

dread feet?

Et quelle puissance et quel art Nouèrent les ligaments de ton

coeur? Dès ses premiers battements, Quels pieds et griffes effrayant?

What the hammer? What the

chain, In what furnace was thy brain? What the anvil? What dread

grasp, Dare its deadly terrors clasp!

Quelle chaîne? Quel marteau? Dans quelle fournaise était ton

cerveau? Quelle enclume? Quel poing ef-

farant Put gripper tes mortelles tran-

ses?

When the stars threw down

their spears And water’d heaven with their

tears: Did he smile his work to see? Did he who made the Lamb ma-

ke thee?

Quand les astres jetant leurs

lances Arrosèrent le ciel de pleurs, Sourit-Il en te contemplant, Lui qui fit l’agneau de douceur?

Tyger Tyger, burning bright, In the forests of the night; What immortal hand or eye, Dare frame thy fearful symme-

try?

Tigre, tigre flamboyant

d’ardeur Dans les forêts de la nuit; Quel pouvoir immortel, inouï, Osa former ta parfaite terreur?

…  …  …

2/2

Author of this English poem is William Blake (∗1757-11-28 – †1827-08-12).

Arg-1048-2125 (2014-06-08 19:43:08)

Pri la poeto vidi la vikipediejon http: // eo. wikipedia. org/ wiki/ William_Blake .

Translation of the English poem “The Tyger ” by William Blake (∗1757-11-28 – †1827-08-12) into franca by Félix Rose.

Arg-1048-2129 (2014-06-09 11:37:35)

Tiu ĉi poem-traduko troviĝas en http: // donh. best. vwh. net/ Esperanto/ Literaturo/ Poezio/ tigro. html .

Translation of the English poem “The Tyger ” by William Blake (∗1757-11-28 – †1827-08-12) into German by Walter A. Aue.

Arg-1048-2126 (2014-06-08 19:55:12)

Tiu ĉi poem-germanigo troviĝas en http: // myweb. dal. ca/ waue/ Trans/ Blake-Tiger. html .

http://www.poezio.net/version?poem-id=1048&version-id=,2125,2129,2126